OpenStudy (toshabay16):

A block of sugar maple wood has a mass of 82.8g and density of 0.69g/cm3 the block is a rectangular prism with a length of 5cm and a width of 4cm what is the height of the maple sugar block?

OpenStudy (erak):

D = m/v <-- density equation 0.69 = 82.8 / v v = 82.8/0.69 v = 120 cm^3 V = lwh 120 = (5)(4)h h = 120/20 h = 6 cm

OpenStudy (toshabay16):

Thanks so much

OpenStudy (erak):

Do you follow along? Ask if you need an explanation

OpenStudy (toshabay16):

I follow, u explained it well

OpenStudy (anitas.):

close it if it's answered ;)